A company's new printer can print checks three times faster than its old printer.
with both printers operating ,the checks can be printed in 6 minutes. how long would it take to print the checks with the new printer only ?

Answer by mananth(16949)

let the old printer's speed be x checks
1/x of the check it does it 1 minute
..
new printer will print at 3x checks
1/3x checks it prints in 1 minute
..
both print in 6 minutes
So 1/6 of the job they do in 1 minute
..
1/x + 1/3x = 1/6
3+1 /3x= 1/6
6*4=3x
3x=24
x= 8 checks by the old printer